OpenStudy (el_arrow):

find an equation of the tangent to the curve at the point corresponding to the given value of the parameter

OpenStudy (el_arrow):

x= t-2sint y=t-2cost

OpenStudy (el_arrow):

0<theta<4pi

OpenStudy (el_arrow):

dx/dt = 1-2cos(t) dy/dt= 1+ 2sin(t) is this right?

OpenStudy (phi):

I would find dx/dt and dy/dt then (dy/dt)/ (dx/dt)= dy/dx = slope of tangent line

OpenStudy (phi):

yes, your dx/dt and dy/dt look ok

OpenStudy (el_arrow):

okay what do i do next?

OpenStudy (el_arrow):

square root both the dy/dt and dx/dt?

OpenStudy (phi):

(dy/dt)/ (dx/dt)= dy/dx = slope of tangent line

OpenStudy (phi):

I am thinking find the equation y= mx + b but if we need to find in terms of t, I would have to think about it

OpenStudy (el_arrow):

okay